# Autodeleveraging: Impossibilities and Optimization

### v1

@@ G.3 The Aggregation Paradox
Appendix H Empirical Methodology
We backtest the mechanisms of §9 on Hyperliquid’s October 10, 2025 autodeleveraging episode. Price shocks are represented by timestamps where the chain records an ADL tranche and are indexed by discrete times $t=1,\dots,T$ and grouped by coin $c$. Each shock carries a realized deficit $D_{t}\geq 0$, the total negative equity of the losers in that liquidation cluster, and an account-level winner capacity vector $w_{t}=(w_{t,1},\dots,w_{t,W_{t}})\in{\mbox{\bf R}}_{+}^{W_{t}}$ measured in USD of available positive PNL. Summing $D_{t}$ across shocks therefore reports the aggregate shortfall that must be socialized during the event.
Controllers either select a scalar severity $\theta_{t}\geq 0$ or a haircut vector $h_{t}\in[0,1]^{\mathcal{W}_{t}}$, generating a per-round budget $B_{t}=\theta_{t}D_{t}$ or $B_{t}=w_{t}^{\top}h_{t}$, respectively. Budgets are always truncated to the realized winner capacity $\sum_{i}w_{t,i}$. The release contains $T=1108$ shocks across 161 coins. Across the positive-deficit shocks this reconstruction captures 304.5304.5M USD of loser shortfall, while the corresponding winner capacity only supports 95.895.8M USD of feasible haircuts, leaving at least 208.6208.6M USD structurally uncovered irrespective of policy.
H.1 Data construction
Event sample.
The data (provided by Hydromancer and SonarX) captured Hyperliquid’s core exchange logs covering 21:15–21:30 UTC and retain every coin with at least one liquidation cluster. There are 998 positive-deficit shocks and 110 shocks with zero deficit (for which policies trivially abstain).
Clustering.
Within each coin, fills are sorted by timestamp and grouped into shocks via a fixed inter-arrival rule: a new cluster begins whenever the gap since the previous fill exceeds $\Delta=5$s. If $t_{k}$ denotes the time of fill $k$, the cluster index obeys $\kappa_{1}=0$ and $\kappa_{k}=\kappa_{k-1}+\mathbf{1}\{t_{k}-t_{k-1}>\Delta\}$. All statistics below are computed at the cluster level.
Account aggregates and capacities.
We aggregate filled trades (‘fills’) by account within a cluster. Positive realized PNL marks winners and contributes to capacity, whereas negative PNL marks losers and contributes to the deficit $D_{t}=\sum_{i\in\mathcal{L}_{t}}(-\mathrm{PNL}_{t,i})$. Capacities are inferred in the order notional $\rightarrow$ $|\texttt{start\_position}\times\texttt{price}|$ $\rightarrow$ $|\texttt{closed\_pnl}|$, providing a conservative haircut limit when full notional quotes are unavailable. Queue priority scores replicate the production ranking by summing realized PNL per account within the cluster.
Production queue replay.
We reconstruct the realized queue budgets $B_{t}^{\mathrm{Q}}$ and haircuts by replaying the published priority order and shaving capacity greedily until the recorded budget is exhausted. This reproduces the overshoot and participation statistics observed on chain and serves as our baseline.
H.2 Losses and objectives
Scalar severity surrogate.
Controllers that only choose $\theta_{t}$ minimize the convex surrogate
$f_{t}(\theta)=-\lambda\,\theta D_{t}+\mu\,(M_{t}-\theta D_{t})_{+}+\nu\,(\theta D_{t}-D_{t})_{+}^{2},$
where $M_{t}=\max_{i}w_{t,i}$ is the largest winner, and $\lambda,\mu,\nu\geq 0$ trade off solvency, shortfall, and overshoot. The subgradient satisfies
$\partial f_{t}(\theta)=-\lambda D_{t}-\mu D_{t}\,\mathbb{1}\{M_{t}>\theta D_{t}\}+2\nu(\theta D_{t}-D_{t})_{+}\,D_{t}.$
We tune $(\lambda,\mu,\nu)$ and the initialization/step-size grid $(\theta_{\text{init}},\eta_{0})$ by coarse search and fix the best setting for all severity controllers.
Vector objective.
When actions are full haircut vectors, we minimize
$\mathcal{L}_{t}(h)={}$ $-\lambda_{\mathrm{vec}}\sum_{i=1}^{W_{t}}w_{t,i}(1-e^{-\beta h_{i}})+\mu_{\mathrm{spr}}\|h\|_{2}^{2}+\phi\,\|h-\bar{\theta}_{t}\mathbf{1}\|_{2}^{2}$
$+\nu_{\mathrm{vec}}\,(w_{t}^{\top}h-D_{t})_{+}^{2}+\gamma_{\mathrm{vec}}\,(D_{t}-w_{t}^{\top}h)_{+}^{2}+\rho_{\mathrm{vec}}\,(D_{t}-w_{t}^{\top}h)_{+},$
with $\bar{\theta}_{t}=\min\{D_{t},\sum_{i}w_{t,i}\}/D_{t}$ when $D_{t}>0$. The weights $(\lambda_{\mathrm{vec}},\beta,\mu_{\mathrm{spr}},\phi,\nu_{\mathrm{vec}},\gamma_{\mathrm{vec}},\rho_{\mathrm{vec}})$ are fixed to the calibrated values shipped with the release. We enforce the budget equality constraint $w_{t}^{\top}h=B_{t}^{\star}$ with $B_{t}^{\star}=\min\{D_{t},\sum_{i}w_{t,i}\}$ via the projection described below.
Revenue loss proxy.
To translate haircut allocations into an expected loss of future fees we use a funding-rate-based proxy. For each winner $i$ in shock $t$ we observe the realized haircut ratio $f_{t,i}=h_{t,i}/w_{t,i}$ and treat it as a churn shock with probability $p_{t,i}=1-\exp(-\beta f_{t,i})$, where $\beta=5$ controls sensitivity. The remaining positive PNL $w_{t,i}$ serves as a conservative proxy for the trader’s deployable notional. We combine this with an estimate of per-dollar daily revenue,
$r^{\text{daily}}_{t}=24\,|f^{\text{fund}}_{t}|+\tau_{\text{turn}}\cdot\text{fee}_{\text{taker}},$
where $|f^{\text{fund}}_{t}|$ is the absolute funding rate observed for the coin (defaulting to $1.5\times 10^{-5}$ when data are sparse), $\tau_{\text{turn}}=1$ assumes one notional turnover per day, and $\text{fee}_{\text{taker}}=4$ bps. Multiplying $r^{\text{daily}}_{t}$ by a two-year horizon $H=730$ days yields the per-dollar lifetime value estimate. The per-shock revenue loss is then
$\text{RevLoss}_{t}=H\,r^{\text{daily}}_{t}\sum_{i}w_{t,i}\,p_{t,i},$
which we aggregate across shocks for each controller.
The exponential hazard ensures that zero haircuts map to zero churn, while large cuts asymptotically drive $p_{t,i}\rightarrow 1$, mirroring the threshold-like exits we observe in the account-level replay data. We set $\beta=5$ so that a 10% haircut implies $\approx 39\%$ churn and a 20% haircut implies $\approx 63\%$ churn, matching the attrition rate measured by comparing pre/post-shock winner mass in the Hyperliquid October 10 dataset [HyperMultiAssetedADL]. Funding rates $f^{\text{fund}}_{t}$ come directly from the “misc events” stream bundled with the same dataset: we aggregate the absolute funding paid per coin over the preceding hour and divide by the coin’s outstanding notional to obtain an hourly rate. When the logs lack funding entries for a coin during the 12-minute window we substitute the cross-sectional median absolute rate of $1.5\times 10^{-5}$ (the BTC/ETH average). This proxy ignores second-order follower dynamics (e.g., strategic re-entry) but provides a consistent way to compare policies on a common revenue scale using only observable quantities.
H.3 Policies
Queue (production).
The production policy orders winners by the realized priority score $q_{t,i}$ and applies the greedy queue allocator with the recorded budget $B_{t}^{\mathrm{Q}}$. Haircuts therefore match the on-chain execution and may overshoot $D_{t}$.
Smart queue.
We retain the same ordering but cap the budget at the feasible mass: $B_{t}^{\mathrm{SQ}}=\min\{D_{t},\sum_{i}w_{t,i}\}$. The greedy queue allocator is applied under this cap, eliminating overshoot while keeping the queue’s concentration.
Exponential backoff.
We impose the per-round PoA cap $\Theta_{t}=\frac{R}{1+R}\,\frac{M_{t}}{D_{t}}$ (uncapped if $R=\infty$) and update
$\theta_{t}=\min\{\theta^{\mathrm{cand}}_{t},\Theta_{t}\},\qquad\theta^{\mathrm{cand}}_{t}=\begin{cases}\theta_{0},&t=1,\\ \max\{\alpha\theta_{t-1},\theta_{0}\},&t\geq 2,\end{cases}$
with $\alpha\in(0,1)$. The resulting budget $B_{t}=\min\{\theta_{t}D_{t},\sum_{i}w_{t,i}\}$ is distributed pro-rata across capacities.
Mirror descent.
We initialize at $\theta_{1}=\theta_{\mathrm{init}}$, compute $g_{t}\in\partial f_{t}(\theta_{t})$, take $\tilde{\theta}_{t+1}=\theta_{t}-\eta_{t}g_{t}$ with $\eta_{t}=\eta_{0}/\sqrt{t}$, perform a short backtracking line-search, and project onto $0,\Theta_{t}]$. Budgets and allocations mirror the backoff policy and differ only through the learning dynamics.
Dynamic2 warm start.
Dynamic2 matches mirror descent but sets $\theta_{1}$ to the levered pro-rata fixed point $\theta_{\mathrm{LPR}}$ computed from the first shock before continuing with the same updates.
Vector mirror descent.
We initialize $h_{t}\equiv 0$ and iterate for $s=1,\dots,S$
[…]
where $\mathcal{C}_{t}=\{h\in[0,1]^{W_{t}}:w_{t}^{\top}h=B_{t}^{\star}\}$. The projection solves
$\min_{h\in[0,1]^{W_{t}}}\tfrac{1}{2}\|h-\tilde{h}\|_{2}^{2}\ \text{s.t.}\ w_{t}^{\top}h=B_{t}^{\star},$
and admits the closed form $h_{i}(\tau)=\big[\tilde{h}_{i}-\tau w_{t,i}\big]_{0}^{1}$ with $\tau$ found by bisection so that $\sum_{i}w_{t,i}h_{i}(\tau)=B_{t}^{\star}$.
H.4 Backtest summary
Coverage.
Aggregate deficits over the event sum to \$304.5304.5M USD. Because winner capacity within most clusters is thin, the feasible budget $B_{t}^{\star}$ totals only 95.895.8M USD, leaving a structural residual of 208.6208.6M USD that no policy can eliminate.
Production queue.
The replayed Hyperliquid queue expends \$704.6704.6M USD of budgets while overshooting deficits by \$630.5630.5M USD and still leaves \$230.3230.3M USD of residual loss. Average winner participation is $31.7\%$, but the queue routinely concentrates haircuts: the mean budget-to-deficit ratio is $9.4\times 10^{3}$ and the largest single haircut reaches \$47.247.2M USD. This misallocation also destroys \$602.8602.8M USD of positive PNL.
Queue variants.
Capping the queue’s budget at feasibility removes the \$630.5630.5M USD overshoot and improves average participation to $60.9\%$, yet it still leaves \$222.1222.1M USD of residual losses and \$113.8113.8M USD of revenue destruction.
Adaptive severity controllers.
[…]
Vector mirror descent.
Optimizing the full haircut vector supplies \$87.587.5M USD of budgets, drives overshoot below $10^{-4}$ USD, and achieves the lowest residual (\$217.0217.0M USD) and second-lowest revenue loss (\$136.9136.9M USD). Winner participation averages $45.2\%$, reflecting the model’s explicit spreading and budget-equality constraint.
Appendix I Formal Proof of the ADL Trilemma
In this appendix, we provide a complete formal statement and proof of the ADL Trilemma (Proposition 2.1). The proof assembles results from Appendices BG, demonstrating that the three-way tension between solvency, fairness, and revenue is a fundamental structural constraint under heavy-tailed shortfalls.
@@ Appendix I Formal Proof of the ADL Trilemma
We work in the large-market limit $n\to\infty$ under the standard assumptions established in §3 and Appendix B.
@@
We impose the following standard assumptions from Appendix B:
Assumption I.1 (Regular Variation).
The right tails of the winner equity distribution $\bar{F}_{+}(x)$ and loser shortfall distribution $\bar{F}_{-}(x)$ are regularly varying with indices $\alpha_{+}>0$ and $\alpha_{-}>0$, respectively:
@@
    Aggregates: $W_{T}/n\xrightarrow{p}\mu_{+}$, $D_{T}/n\xrightarrow{p}\mu_{-}$, and total fees $\Phi_{T}/n\xrightarrow{p}\mu_{\Phi}$ for constants $\mu_{\pm},\mu_{\Phi}\in(0,\infty)$.
@@
    Extremes: The maximum winner equity $\omega_{T}$ and maximum loser shortfall $\Delta_{T}$ satisfy
    $\frac{\omega_{T}}{b_{k_{n}}^{+}}\xrightarrow{p}c_{+},\qquad\frac{\Delta_{T}}{b_{m_{n}}^{-}}\xrightarrow{p}c_{-},$
@@
Assumption I.3 (Structural Deficit Regime).
@@
Definition I.5 (Fairness / Bounded Moral Hazard).
@@
    $c_{\text{lo}}\;\leq\;\mathsf{PTSR}_{T}(\mathcal{P}_{n},\pi_{n}):=\mathop{\bf E{}}\left[\frac{\omega_{T}^{\pi_{n}}}{D_{T}^{\pi_{n}}}\right]\;\leq\;c_{\text{hi}};$
@@
    $c^{\prime}_{\text{lo}}\;\leq\;\mathsf{PMR}_{T}(\mathcal{P}_{n},\pi_{n}):=\mathop{\bf E{}}\left[\frac{\omega_{T}^{\pi_{n}}}{\Delta_{T}^{\pi_{n}}}\right]\;\leq\;c^{\prime}_{\text{hi}}.$
These bounds ensure the top winner’s residual equity remains proportional to the deficit scale. For brevity we write $\mathsf{PTSR}_{T}(\pi_{n})$ (and $\mathsf{PMR}_{T}(\pi_{n})$) whenever the dependence on $\mathcal{P}_{n}$ is clear from context.
Definition I.6 (Revenue Preservation).
@@
Theorem I.7 (ADL Trilemma).
Let $(\mathcal{P}_{n})_{n\geq 1}$ be a sequence of perpetuals exchanges satisfying Assumptions I.1I.3. For any static ADL policy family $(\pi_{n})$ with severity sequence $(\theta_{n})$, at most two of the three conditions (S), (F), and (R) can hold simultaneously.
@@
I.1 Proof of the Trilemma
@@
Lemma I.8 (Solvency-Revenue Identity).
@@
The proof then proceeds by analyzing this identity under different constraint combinations.
Proof of Implication (I): (S) $\wedge$ (F) $\Rightarrow$ $\neg$(R).
@@
Suppose both (S) and (F) hold.
1. Fairness Implication: By (F), $\theta_{n}=O(b_{n}/n)$ (Theorem B.3). Since $b_{n}=o(n)$, the haircut budget scales as:
$B_{T}=\theta_{n}D_{T}\asymp\frac{b_{n}}{n}\cdot n=b_{n}=o(n).$
Thus, $B_{T}/n\xrightarrow{p}0$.
2. Solvency Implication: By (S), residual risk is negligible, so $R_{T}/n\xrightarrow{p}0$.
3. Identity Analysis: Using the Solvency-Revenue Identity (divided by $n$):
$\frac{\mathrm{LTV}_{T}}{n}\;\leq\;\frac{\Phi_{T}}{n}+\frac{B_{T}}{n}+\frac{R_{T}}{n}-\frac{D_{T}}{n}+\frac{o(n)}{n}.$
Taking limits $n\to\infty$:
$\lim_{n\rightarrow\infty}\frac{\mathrm{LTV}_{T}}{n}\;\leq\;\mu_{\Phi}+0+0-\mu_{-}+0.$
By Assumption I.3, $\mu_{-}>\mu_{\Phi}$. Thus:
$\lim_{n\rightarrow\infty}\frac{\mathrm{LTV}_{T}}{n}\;\leq\;\mu_{\Phi}-\mu_{-}\;<\;0.$
This inequality implies a contradiction for any solvent, self-sustaining exchange. Since $\mathrm{LTV}_{T}$ represents retained earnings, a strictly negative limit implies the exchange requires external subsidies to cover the structural deficit. In the absence of external capital, the exchange must either halt or become insolvent. Consequently, $\mathrm{LTV}_{T}(\pi_{n})$ cannot satisfy the revenue condition (R), which requires $\mathrm{LTV}_{T}$ to be a positive fraction of $\Phi_{T}$. ∎
Proof of Implication (II): (S) $\wedge$ (R) $\Rightarrow$ $\neg$(F).
Proof.
Suppose both (S) and (R) hold.
1. Revenue Implication: By (R), $\mathrm{LTV}_{T}\geq c_{R}\Phi_{T}$. This implies diversions are limited: $\mathcal{D}_{T}=\Phi_{T}-\mathrm{LTV}_{T}\leq(1-c_{R})\Phi_{T}$.
2. Solvency Implication: By (S), $R_{T}\approx 0$. From $D_{T}\leq B_{T}+\mathcal{D}_{T}+R_{T}$, we have:
$B_{T}\;\geq\;D_{T}-\mathcal{D}_{T}\;\geq\;D_{T}-(1-c_{R})\Phi_{T}.$
3. Fairness Violation: Dividing by $n$ and taking limits:
$\lim_{n\rightarrow\infty}\frac{B_{T}}{n}\;\geq\;\mu_{-}-(1-c_{R})\mu_{\Phi}.$
Since $\mu_{-}>\mu_{\Phi}$ (Assumption I.3) and $c_{R}\in(0,1]$, the RHS is strictly positive:
$\mu_{-}-(1-c_{R})\mu_{\Phi}\;>\;\mu_{\Phi}-(1-c_{R})\mu_{\Phi}\;=\;c_{R}\mu_{\Phi}\;>\;0.$
Thus, $B_{T}=\Theta(n)$. Since $B_{T}=\theta_{n}D_{T}$ and $D_{T}=\Theta(n)$, this implies $\theta_{n}=\Theta(1)$.
With $\theta_{n}=\Theta(1)$, the load $\kappa_{n}=\theta_{n}n/b_{n}\to\infty$ (since $b_{n}=o(n)$). By Theorem B.3, $\mathsf{PTSR}_{T}\asymp 1/\kappa_{n}\to 0$. This violates (F), which requires $\mathsf{PTSR}_{T}=\Theta(1)$. ∎
Proof of Implication (III): (F) $\wedge$ (R) $\Rightarrow$ $\neg$(S).
Proof.
Suppose both (F) and (R) hold.
1. Fairness Implication: As shown in (I), (F) implies $B_{T}=o(n)$.
2. Revenue Implication: As shown in (II), (R) implies $\mathcal{D}_{T}\leq(1-c_{R})\Phi_{T}$.
3. Solvency Violation: From the Solvency-Revenue Identity: $R_{T}\geq D_{T}-B_{T}-\mathcal{D}_{T}$. Dividing by $n$ and taking limits:
$\lim_{n\rightarrow\infty}\frac{R_{T}}{n}\;\geq\;\mu_{-}-0-(1-c_{R})\mu_{\Phi}.$
By Assumption I.3 ($\mu_{-}>\mu_{\Phi}$), we have:
$\mu_{-}-(1-c_{R})\mu_{\Phi}\;>\;\mu_{\Phi}-(1-c_{R})\mu_{\Phi}\;=\;c_{R}\mu_{\Phi}\;>\;0.$
Thus, $R_{T}$ scales as $\Theta(n)$. This implies $\mathop{\bf Prob}[R_{T}>0]\to 1$ and $\sum R_{t}\to\infty$, violating (S). ∎
@@ 3. Fairness Violation: Dividing by $n$ and taking limits:
Proposition I.9 (Attainability of Two Desiderata).
Under Assumptions I.1I.3:
@@
I.2 Connection to Classical Impossibility Results
@@
  • Arrow’s Impossibility Theorem: No voting rule satisfies Pareto efficiency, independence of irrelevant alternatives, and non-dictatorship simultaneously [Arrow1951].
  • Mundell-Fleming Trilemma: In international finance, a country cannot simultaneously maintain a fixed exchange rate, free capital movement, and independent monetary policy [Mundell1963, Fleming1962].
  • CAP Theorem: In distributed systems, a database cannot provide consistency, availability, and partition tolerance simultaneously [Brewer2000, GilbertLynch2002].
  • Credibility Trilemma: Single-item auctions cannot be simultaneously optimal, strategy-proof, and credible, forcing designers to sacrifice at least one desideratum [AkbarpourLi2020].
@@
    Quantitative Arrow’s Theorem: Recent results in quantitative social choice show that while Arrow’s impossibility holds in the worst case, the probability of paradoxical outcomes (like intransitivity) can be small for many natural distributions of preferences [MosselNeemanTamuz2014]. Analogously, our ADL trilemma bounds hold with high probability under heavy-tailed distributions, but dynamic policies (like Stackelberg controllers) can minimize the frequency of trilemma-binding events, achieving a “quantitative” relaxation.

### v3

@@ H.3 The Aggregation Paradox
Appendix I Empirical Methodology
This appendix documents the OSS reproduction pipeline underlying §9. The goal is to make the empirical objects file-backed (every number in §9 is computed from on-disk artifacts under OSS/out/) and to keep the two-space hygiene explicit: production ADL executes in contract space, while we measure impacts in wealth space.
I.1 Data construction
Data sources.
We use HyperReplay’s ground-truth fill and misc event streams together with clearinghouse snapshots to replay winner account states. We also use the canonical REALTIME table adl_detailed_analysis_REALTIME.csv to extract winner sets and to form equity/PNL capacity proxies. The OSS code resolves upstream paths in OSS/src/oss_adl/paths.py and writes derived artifacts under OSS/out/.
Wave construction (global time waves).
We define global waves by gap clustering on ADL-fill timestamps: sort ADL-fill rows by time, start a new wave when the inter-fill gap exceeds $\texttt{gap\_ms}=5000$ms, and define wave $t$ as $t_{\mathrm{start}}(t),t_{\mathrm{end}}(t)]$. Global (not per-coin) waves avoid double counting a single solvency episode across multiple markets.
Loser deficit $D_{t}$.
For each wave, the realized loser deficit is
$D_{t}=\sum_{j\in\mathrm{losers}(t)}(-e_{t}(j))_{+},$
computed from loser-side liquidation equity fields. Concretely, for each liquidated_user we take the minimum observed liquidated_total_equity within the wave and sum the negative minima across users.
Needed budget $B_{t}^{\mathrm{needed}}$.
For each ADL fill $k$, we parse the liquidation mark $p_{k}^{\mathrm{mark}}$ and execution price $p_{k}^{\mathrm{exec}}$ from the raw fill stream (fields markPx and px), together with size $q_{k}$, and set
$\mathrm{needed}_{k}:=\lvert p_{k}^{\mathrm{mark}}-p_{k}^{\mathrm{exec}}\rvert\cdot\lvert q_{k}\rvert,\qquad B_{t}^{\mathrm{needed}}:=\sum_{k\in\text{fills in wave }t}\mathrm{needed}_{k}.$
This is an instantaneous bankruptcy-gap proxy: it does not include opportunity cost after the wave ends.
Two-pass replay and production wealth removed $H_{t}^{\mathrm{prod}}(\Delta)$.
We replay the realized event stream twice for the winner set in each wave: an ADL-on pass (apply all fills) and a no-ADL pass (skip ADL fills in state updates while keeping the realized price path). We evaluate equities at $t_{\mathrm{eval}}=t_{\mathrm{end}}+\Delta$ (after wave end we update only the price path) and compute
$H_{t}^{\mathrm{prod}}(\Delta)=\sum_{u\in W(t)}\bigl(e_{t,\mathrm{end}}^{\mathrm{no\text{-}ADL}}(u;\Delta)-e_{t,\mathrm{end}}^{\mathrm{ADL}}(u;\Delta)\bigr)_{+}.$
The no-ADL counterfactual is not identifiable from public data without an assumption; we use the standard “hold the realized price path fixed, remove the ADL state update” counterfactual.
Production overshoot vs needed.
We report
$O_{t}(\Delta):=H_{t}^{\mathrm{prod}}(\Delta)-B_{t}^{\mathrm{needed}},\qquad O(\Delta):=\sum_{t}O_{t}(\Delta).$
A horizon sweep varies $\Delta$ (scenario parameter) to expose a short-horizon opportunity-cost channel. The per-horizon totals are written to OSS/out/eval_horizon_sweep_gap_ms=5000.csv and summarized in OSS/out/overshoot_robustness.json.
I.2 Benchmark allocations targeting $B_{t}^{\mathrm{needed}}$
To make “excessive relative to alternatives” concrete, we compare production to transparent benchmark allocations that target $B_{t}^{\mathrm{needed}}$ wave-by-wave under a profits-only capacity constraint. For each wave we build a per-user capacity proxy
$c_{u}:=\min\{U(u),E(u)\},$
where $U(u)$ is positive unrealized PNL and $E(u)$ is positive equity, both taken from the canonical REALTIME table within the wave window. Benchmarks differ mainly in how they handle contract discreteness:
  • Wealth pro-rata (continuous): capped pro-rata in wealth-space USD over $c_{u}$.
  • Vector mirror descent (vector-md): a projection-based allocator that produces $h=x\odot c$ with $c^{\top}x=B_{t}^{\mathrm{needed}}$ and $x\in[0,1]^{n}$.
  • Contract pro-rata: standard exchange-style integer allocation proportional to position size.
  • Min-max ILP: MIP solver minimizing maximum haircut percentage.
These outputs are written to OSS/out/policy_per_wave_metrics.csv and plotted in Figures 56.
I.3 Diagnostic decomposition (not the §9 overshoot metric)
This subsection is a diagnostic for interpreting execution timing and for answering common “markout” questions; it is not the overshoot metric used in §9. From the ADLed user’s perspective, with signed quantity $Q$, execution price $p_{\text{exec}}$, nearest mark at execution time $p_{\text{mark}}(t)$, and mark at horizon $p_{\text{mark}}(t+\Delta)$, total markout decomposes as
$Q\cdot(p_{\text{mark}}(t+\Delta)-p_{\text{exec}})=Q\cdot(p_{\text{mark}}(t)-p_{\text{exec}})+Q\cdot(p_{\text{mark}}(t+\Delta)-p_{\text{mark}}(t)).$
Because this diagnostic combines executions with time-indexed marks and the public-data replay lacks full clearing/settlement observability, its aggregate need not satisfy a strict two-party zero-sum identity.
Revenue proxy.
To translate haircut allocations into an expected loss of future fees, we use a minimal churn proxy. For a winner $u$ in wave $t$ with USD haircut $h_{t,u}$ and capacity proxy $c_{u}$, define the churn probability
$p_{t,u}=1-\exp\!\left(-\beta\,h_{t,u}/c_{u}\right),$
and use this to scale down a simple fee proxy based on notional and per-trade fee rates. This proxy is intended for within-paper comparisons under a common set of assumptions; it is not an identification claim about venue-level revenue.
Appendix J Formal Proof of the ADL Trilemma
In this appendix, we provide a complete formal statement and proof of the ADL Trilemma (Proposition 2.1). The proof assembles results from Appendices CH, demonstrating that the three-way tension between solvency, fairness, and revenue is a fundamental structural constraint under heavy-tailed shortfalls.
@@ I.3 Diagnostic decomposition (not the §9 overshoot metric)
We work in the large-market limit $n\to\infty$ under the standard assumptions established in §3 and Appendix C.
@@
We impose the following standard assumptions from Appendix C:
Assumption J.1 (Regular Variation).
The right tails of the winner endowment distribution $\bar{F}_{+}(x)$ (under PNL-only, the distribution of positive PNL) and loser shortfall distribution $\bar{F}_{-}(x)$ are regularly varying with indices $\alpha_{+}>0$ and $\alpha_{-}>0$, respectively:
@@
    Aggregates: $U_{T}/n\xrightarrow{p}\mu_{+}$ (total endowment capacity), $D_{T}/n\xrightarrow{p}\mu_{-}$ (total deficit), and total fees $\Phi_{T}/n\xrightarrow{p}\mu_{\Phi}$ for constants $\mu_{\pm},\mu_{\Phi}\in(0,\infty)$.
@@
    Extremes: The maximum winner endowment $\upsilon_{T}$ (under PNL-only, maximum positive PNL) and maximum loser shortfall $\Delta_{T}$ satisfy
    $\frac{\upsilon_{T}}{b_{k_{n}}^{+}}\xrightarrow{p}c_{+},\qquad\frac{\Delta_{T}}{b_{m_{n}}^{-}}\xrightarrow{p}c_{-},$
@@
Assumption J.3 (Structural Deficit Regime).
@@
Definition J.5 (Fairness / Bounded Moral Hazard).
@@
    $c_{\text{lo}}\;\leq\;\mathsf{PTSR}_{T}(\mathcal{P}_{n},\pi_{n}):=\mathop{\bf E{}}\left[\frac{\upsilon_{T}^{\pi_{n}}}{D_{T}^{\pi_{n}}}\right]\;\leq\;c_{\text{hi}},$
    where $\upsilon_{T}^{\pi_{n}}$ is the maximum post-ADL endowment (under PNL-only, maximum post-ADL positive PNL);
@@
    $c^{\prime}_{\text{lo}}\;\leq\;\mathsf{PMR}_{T}(\mathcal{P}_{n},\pi_{n}):=\mathop{\bf E{}}\left[\frac{\upsilon_{T}^{\pi_{n}}}{\Delta_{T}^{\pi_{n}}}\right]\;\leq\;c^{\prime}_{\text{hi}}.$
These bounds ensure the top winner’s residual endowment (profit capacity) remains proportional to the deficit scale. For brevity we write $\mathsf{PTSR}_{T}(\pi_{n})$ (and $\mathsf{PMR}_{T}(\pi_{n})$) whenever the dependence on $\mathcal{P}_{n}$ is clear from context.
Definition J.6 (Revenue Preservation).
@@
Theorem J.7 (ADL Trilemma).
Let $(\mathcal{P}_{n})_{n\geq 1}$ be a sequence of perpetuals exchanges satisfying Assumptions J.1J.3. For any static ADL policy family $(\pi_{n})$ with severity sequence $(\theta_{n})$, at most two of the three conditions (S), (F), and (R) can hold simultaneously.
@@
Remark J.8 (Interpretation of $\mu_{\Phi}$).
In applications, $\mu_{\Phi}$ can be interpreted as the largest fee diversion rate that remains compatible with non-declining long-run venue value (e.g., via an LTV sensitivity constraint). This is an interpretation of the regime boundary, not a change to the formal assumption used in the proof.
Remark J.9 (Theorem vs observability).
Theorem J.7 is proved in the policy model and does not assume any particular empirical observability of execution/settlement cashflows. Empirical tests necessarily operate under an observation model and should be interpreted as measurements of the paper’s defined objects, not as complete ledger identities.
J.1 Proof of the Trilemma
@@
Lemma J.10 (Solvency-Revenue Identity).
@@
By definition of residual insolvency, any uncovered shortfall after applying haircuts and diverted fees must appear as $R_{T}$:
$R_{T}\;\geq\;D_{T}-B_{T}-\mathcal{D}_{T}-K_{0},$
which rearranges to $D_{T}\leq B_{T}+\mathcal{D}_{T}+K_{0}+R_{T}$, establishing the first inequality. Substituting $\mathcal{D}_{T}=\Phi_{T}-\mathrm{LTV}_{T}$ and using $K_{0}=o(n)$ yields the asymptotic bound. ∎
Proof of Theorem J.7.
(S) and (F) $\Rightarrow$ $\neg$(R). By (F) and Theorem C.3, $\theta_{n}=O(b_{n}/n)$, so $B_{T}=o(n)$; (S) gives $R_{T}=o_{p}(n)$. Lemma J.10 implies
$\frac{\mathrm{LTV}_{T}}{n}\;\leq\;\mu_{\Phi}+o(1)-\mu_{-}\;<\;0$
using Assumption J.3, contradicting (R).
(S) and (R) $\Rightarrow$ $\neg$(F). Revenue bounds diversions, so solvency requires $B_{T}\gtrsim n$, hence $\theta_{n}=\Theta(1)$. Then $\kappa_{n}=\theta_{n}n/b_{n}\to\infty$ (since $b_{n}=o(n)$); Theorem C.3 gives $\mathsf{PTSR}_{T}\to 0$, violating (F).
(F) and (R) $\Rightarrow$ $\neg$(S). Fairness again yields $B_{T}=o(n)$; revenue caps $\mathcal{D}_{T}$. Lemma J.10 gives
$\frac{R_{T}}{n}\;\geq\;\mu_{-}-(1-c_{R})\mu_{\Phi}-o(1)\;>\;0,$
so $\mathop{\bf Prob}[R_{T}>0]\to 1$, violating (S). ∎
@@
Proposition J.11 (Attainability of Two Desiderata).
Under Assumptions J.1J.3:
@@
J.2 Connection to Classical Impossibility Results
@@
  • Arrow’s Impossibility Theorem: No voting rule satisfies Pareto efficiency, independence of irrelevant alternatives, and non-dictatorship simultaneously [13].
  • Mundell-Fleming Trilemma: In international finance, a country cannot simultaneously maintain a fixed exchange rate, free capital movement, and independent monetary policy [88, 55].
  • CAP Theorem: In distributed systems, a database cannot provide consistency, availability, and partition tolerance simultaneously [30, 59].
  • Credibility Trilemma: Single-item auctions cannot be simultaneously optimal, strategy-proof, and credible, forcing designers to sacrifice at least one desideratum [7].
@@
    Quantitative Arrow’s Theorem: Recent results in quantitative social choice show that while Arrow’s impossibility holds in the worst case, the probability of paradoxical outcomes (like intransitivity) can be small for many natural distributions of preferences [87]. Analogously, our ADL trilemma bounds hold with high probability under heavy-tailed distributions, but dynamic policies (like Stackelberg controllers) can minimize the frequency of trilemma-binding events, achieving a “quantitative” relaxation.